# Unlocking the ‘Why’ of Buying: Introducing a New Dataset and Benchmark for Purchase Reason and Post-Purchase Experience

### v1

@@ Abstract
In our work, we propose a novel purchase reason explanation task. To this end, we introduce an LLM-based approach to generate a dataset that consists of textual explanations of why real users make certain purchase decisions. We induce LLMs to explicitly distinguish between the reasons behind purchasing a product and the experience after the purchase in a user review. An automated, LLM-driven evaluation, as well as a small scale human evaluation, confirms the effectiveness of our approach to obtaining high-quality, personalized explanations. We benchmark this dataset on two personalized explanation generation tasks. We release the code and prompts to spur further research [^1] The URL will be provided after the anonymity period..
@@
[…]
@@ 1 Introduction
Providing user-understandable explanations to justify recommendations could enhance the effectiveness, persuasiveness, and user satisfaction of recommendation Zhang and Chen (2020). This has been confirmed by both user studies Zanker (2012) and many real products such as Microsoft Office 365 Xu et al. (2020), JD.com (an e-commerce website) Zhang et al. (2014a) and Spotify McInerney et al. (2018). Explanations can be presented in various styles, such as a piece of text, a relevant user or item, a radar chart, an image or a set of reasoning rule Tintarev and Masthoff (2015). With the advancement of natural language generation techniques, a short narrative (e.g., a sentence) becomes one of the most popular explanation style, which is the focus of our work.
To pave the way of building a more effective recommendation system, we need datasets that explain the reasoning behind user choices. Existing works leverage user reviews to mine explanations. One work Li et al. (2021a) extracts the most commonly occurred (near) duplicate sentences across reviews, resulting in short and generic comments about the items. For instance, “Excellent movie” is the top 1 extracted explanation in Amazon Movies & TV review dataset. Others extract review sentences Geng et al. (2022) or segments Ni et al. (2019) that mention one or more pre-generated item features/aspects. An example explanation is “The quality of the material is great” where “quality” and “material” are two features of the item Ni et al. (2019).
@@
In this work, we introduce a novel purchase reason explanation task. Our approach leverages a language model (LLM) to generate a dataset from user reviews, capturing purchase reasons and in the meantime, generating a highly relatable, personalized post-purchase experience. We propose four dimensions to measure the data quality and validate the effectiveness of our approach through automated LLM evaluator and small-scale human feedback. The resulting dataset is a high-quality, personalized set of explanations. With this dataset, we can develop models to generate explanations from two aspects, the relevance to the user need (i.e., purchase reason) and the preference for the particular item (i.e., post-purchase experience). We benchmark the dataset against these two generation tasks with an LLM and explore different user and item representations.
@@
  • To spur further research, we release the code and all the prompts for generating the dataset and benchmarking.
@@ 2 Related Work
In this section, we first review prior work on dataset construction for personalized explanations and explanation generation methods in the context of explainable recommender systems. We then discuss the use of large language models (LLMs) for text generation and evaluation.
@@ 2.1 Personalized Explanation Dataset
Constructing a high quality personalized explanation dataset is the key to build an explainable model for personalized recommendation. Prior work extracts various information from user reviews, such as using the entire reviews Chen et al. (2018); Li et al. (2017), sentences Chen et al. (2019); Wang et al. (2018b), aspect-specific sentences Geng et al. (2022), or elementary discourse units Ni et al. (2019). They discard segments that are too personal (containing first-person or third-person pronouns) or too short/long. Li et al. (2021a) extract commonly occurred near-duplicate sentences as explanations, which are often short and generic (e.g., “Excellent movie”). The above approaches capture customer satisfaction rather than purchase motives because they do not distinguish pre- from post-purchase experiences.
Li et al. (2017) used Tips [^2] https://www.yelp-support.com/article/What-are-tips?l=en_US, a concise form of reviews, as justifications for choosing certain businesses. But Tips have low availability (e.g., Yelp mobile). To our knowledge, we are the first to extract purchase reasons directly from regular reviews for creating personalized explanation datasets.
@@ 2.2 Personalized Explanation Generation
Explainable recommendation Zhang and Chen (2020) has drawn considerable research attention. We limit our discussion to personalized, textual explanation generation, which aims to justify why the recommended item might match a user’s interest. Ranking based generation Li et al. (2021a, 2023a) selects explanations from a pre-generated candidate pool, where explanations could be represented as IDs. Template-based explanation Zhang et al. (2014b); Wang et al. (2018a) selects candidate words to fill into a template sentence. Both approaches tend to generate generic explanations, lacking language flexibility and personalization.
Natural language generation (NLG) is widely used to generate free-text explanations. Early studies fine-tuned seq2seq models like LSTM Costa et al. (2018), GRU Li et al. (2017); Ni et al. (2019), Transformer Li et al. (2021b) and T5 Liu et al. (2023) or GPT-2 Li et al. (2023b). Limited by the model capability, however, the generated explanations are often generic and not fluent Liu et al. (2023). Recent work with powerful LLMs (e.g., ChatGPT Liu et al. (2023)) in zero or few-shot setups significantly improved quality.
A common challenge for NLG based methods is representing users and items. Most studies use textual descriptions of users (e.g., reviews) and items (e.g., title). User and item IDs are also considered Liu et al. (2023); Li et al. (2023b), which are helpful for capturing interactions while limiting generalization to new IDs.
A shared challenge of all these approaches is how to systematically evaluate the generated personalized explanation, which we address by introducing a new dataset and benchmark.
2.3 Text Generation and Evaluation With LLMs
LLMs have shown impressive abilities in generation tasks Zhao et al. (2023), matching commercial translation tools Jiao et al. (2023) and human writers Zhang et al. (2024). However, they can produce hallucinations Huang et al. (2023). Readers can refer to a few recent survey papers for comprehensive discussions Zhang et al. (2023); Huang et al. (2023).
LLMs have been widely used to evaluate the quality of generated text Zhao et al. (2023) across various domains such as summarization Wang et al. (2023a), translation Kocmi and Federmann (2023) and personalized text generation Wang et al. (2023b), often achieving state-of-the-art or competitive correlation with human judgments.
Our work is related to this general line of research as we use LLMs to generate purchase reasons from user reviews and to evaluate the quality of the extracted explanation.
@@ 3 Generating the Explanation Dataset
As we discussed in Section 2.1, prior work that extracts potential explanations from user reviews has several key limitations. Firstly, most of existing work heavily relies on the user’s sentiment on the item rather than their initial motives. As the reviews are generally written after purchasing the item, it is more common that these sentiments stem from user experience that is post-purchase, rather than the original reasons that motivate the users to purchase. Consequentially, the extracted information may well explain the user’s rating of the item, but it might not be a good explanation of their purchase decision. Secondly, the prior generated explanations tend to be generic and non-personal. For instance,  Li et al. (2021a) extracted the most frequently occurring near-duplicate sentences as explanations, and  Ni et al. (2019) discarded text segments that are personal (containing first-person or third-person pronouns). This is not hard to understand: while the assessments of an item are often shared across a majority of users, the reasons for making a purchase can be much more personal.
@@
To address these limitations, we explicitly distinguish the purchase reasons from post-purchase experience. These two concepts have strong roots in the business and marketing literature, where purchase reason is related to purchase intent (or intention) Chang and Wildt (1994) and purchase motive (or motivation) Sirgy (1985), while post-purchase experience is related to customer (or product) satisfaction Oliver and Swan (1989); Richins and Bloch (1991). In particular, purchase intent refers to the likelihood that a user will buy a product, which is influenced by a number of factors such as personal needs, customer preferences, brand equity, trust, perceived value of the product, etc. Cobb-Walgren et al. (1995); Bleize and Antheunis (2019) In our work, we do not further distinguish these factors and in general refer to them as purchase reasons. On the other hand, post-purchase experience encodes a variety of factors that influences post-purchase customer satisfaction, which is highly related to the user’s attitude towards the product or business and may further influences their future purchasing behavior Anderson et al. (1979). In our work, the practical difference between purchase reason and post-purchase experience lies in whether the conveyed information is before the user purchases the item or after. The former explains the user’s decision to purchase the item, and the latter explains the decision of rating the item.
@@
As LLMs (eg, ChatGPT [^3] https://openai.com/chatgpt, GPT-4 OpenAI (2023), Gemini Gemini Team (2023)) have demonstrated strong performance in text generation, we propose to utilize LLMs to generate both types of explanations based on user reviews, rather than adopting extractive methods by the prior works. We carefully devise strategies to improve generation quality and combat hallucinations. Moreover, we use LLMs as an auto-rater to assess the generated explanations from multiple aspects and further improve data quality. To be specific, we demonstrate our solution with Gemini Ultra and Amazon product review 5-core dataset  [^4] https://nijianmo.github.io/amazon/index.htmlNi et al. (2019). We detail our end-to-end solution in the following.
3.1 Generating Explanations using LLM
We formally define the goal: given a product and an associated user review, asking an LLM (Gemini Ultra in our case) to generate the user’s purchase reason and post-purchase experience for this product. We start with generating the two explanations as two separate tasks and identify a few issues with the initial exploration. Firstly, we find an LLM does not have a clear distinction between purchase reasons and post-purchase experience. The two generated explanations have overlaps or appear in the wrong explanation group. Secondly, we notice that users do not always explicitly express their purchase reasons in the review. In those cases, the model either infers the potential purchase reasons from the product information (e.g., some highlighted features by sellers) or generates them by hallucination. We devise the following strategies to address the issues and list the final prompt in Appendix (Table 12).
  • We ask the model to perform the task of purchase reason and post-purchase experience generation simultaneously in one prompt. This forces the model to draw a clear boundary between the two types of explanations.
  • To alleviate hallucination, we break down purchase reasons into explicit (reasons the user explicitly mentioned in the review) and implicit (reasons that can be inferred from the review and the product information). We allow the model to leave any of them as empty when no enough information is provided.
@@ 3.1 Generating Explanations using LLM
  • We require the model to provide supporting evidence for its generated purchase reasons. This helps combat hallucination and promotes more accurate and concise explanations.
We also experiment with including few-shot examples in in-context learning for the explanation generation task. We do not find significant performance gains by including few-shot examples.
3.2 Rating Explanations using LLM
@@
Prior work has demonstrated a great success of automatic evaluation of text generation with LLMs (detailed in Section 2.3). Inspired by this, we explore the usage of LLMs in assessing the quality of our explanation dataset. In particular, we propose to measure the quality of purchase reasons and post-purchase experience from the following four dimensions. We conduct two separate evaluations for the two types of explanations (see Table 13 and Table 14 in Appendix for the specific prompts). We further leverage the four evaluation results to filter out noises, aiming at improving the dataset quality.
@@
  • Correctness: if the model correctly identifies the explanation type. We observe that the model sometimes confuse post-purchase experience as purchase reason, but rarely the opposite.
  • Completeness: if the explanation covers all relevant aspects present in the product and the review.
  • Personalization: if the explanation contains adequate personal context and is not too generic. It is common that some reviews are generic. This is not a quality metric but a tool to characterize the personalization level.
@@ 4 Dataset Evaluation and Analysis
As mentioned, we demonstrate our LLM based solution with Gemini Ultra and Amazon product review 5-core dataset Ni et al. (2019). The full 5-core dataset consists of reviews from all users and items that have at least 5 reviews, resulting in 75 million reviews in total. Considering the cost of using LLMs, we randomly sample 10K reviews for experiments. We apply the LLM generator to construct the explanation dataset and utilize the LLM auto-rater to judge the data quality. In the following, we first validate the effectiveness of our LLM auto-rater and then describe the characteristics of the generated explanation dataset.
@@ 4.1 Effectiveness of LLM Auto-rater
To evaluate the effectiveness of the LLM auto-rater, we conduct a small scale human annotation with four annotators by following the guidelines as the auto-rater uses. As a pilot study, we ask human annotators to label explanations for 20 reviews. We find there is a high agreement among annotators: they have perfect consensus on hallucination and personalization dimension, and pick up a common answer for 90% of explanations in terms of correctness and completeness. We further ask annotators to discuss the conflicting answers to reach a consistent interpretation of the guidelines. In the later formal annotation, we work on a larger set of 100 randomly sampled reviews, and each explanation is annotated by one person.
As detailed in Table 2, we find LLM auto-rater achieves a strong correlation with human judgement in hallucination (agree on 96% of reviews for purchase reason and 98% for post-purchase experience) and correctness (92% of reviews). However, they disagree more on completeness measurement, only reaching an agreement on 84% and 80% of reviews for purchase reason and post-purchase experience, respectively. The main reason is that the LLM auto-rater is good at capturing subtle information presented in the review and the product, and thus is more strict than human. For instance, human label purchase reasons for 91% of reviews as complete while auto-rater is 77% only. The personalization annotation, aiming at characterizing the explanations, is more challenging due to its subjective nature. The agreement between the LLM auto-rater and human rater is 73% and 80% for purchase reason and post-purchase experience, respectively. The LLM auto-rater labels more explanations as personal than human, indicating human are more conservative than an LLM in judging personalization.
Table 2: The percentage of examples that LLM auto-rater and human annotator agree. “ Hall.”/“Compl.” denote hallucination and completeness, respectively.
Hall. Correct Compl. Personal
Reason 96% 92% 84% 73%
Experience 98% N.A. 80% 80%
@@ 4.2 Quality of Generated Explanations
We show the evaluation results of the generated explanations by the LLM auto-rater in Table 3. To understand how the quality varies across products, we show the break down results for top five product categories. Overall, the LLM generator performs very well in combating hallucination with a very low hallucination rate of 0.49% for purchase reason and 0.21% for post-purchase experience. Moreover, the LLM generator has a clear distinction about the two explanations, only confusing post-purchase experience as purchase reason in 0.61% of cases.
As expected, the explanation complete rate is low (77.23% for purchase reason and 77.23% for post-purchase experience). As discussed in Section 4.1, the LLM auto-rater is very strict in completeness evaluation, aiming at capturing all the subtle explanations. This suggests that the actual completeness rate could be higher than the LLM auto-rater detected.
Within the five product categories, we find that “Electronics” has the highest purchase reason complete rate (81.85%), while “Clothing, Shoes & Jewelry” is the lowest (71.49%). We hypothesize the discrepancy might due to the different purchase and review behaviors for the two categories. Unlike electronics buyers, who often explicitly explain their purchase motivations in reviews, those buying clothes, shoes, or jewelry are less likely to do so. This means an LLM must infer these motivations from more subtle cues within the reviews and is more prone to neglect some reasons. For post-purchase experience, we also observe that “Electronics” has the highest complete rate (72.51%) than other product categories. The possible reason might be that “Electronics” reviews tend to focus on the product features, and use more direct languages, which is easy for an LLM to capture all post-purchase experience.
Table 3: The auto-rater results on purchase reason. Fashion denotes Clothing, Shoes & Jewelry.
All Books Fashion Home & Kitchen Electronics Sports & Outdoors
@@
Hallucination Rate 0.49% 0.38% 0.62% 0.37% 0.44% 0.31%
Correct Rate 99.39% 99.46% 99.52% 99.08% 99.56% 99.07%
Complete Rate 77.23% 73.30% 71.49% 77.29% 81.85% 80.00%
Personal Rate 70.01% 78.89% 68.32% 67.12% 61.50% 66.05%
@@
Hallucination Rate 0.21% 0.22% 0.07% 0.19% 0.11% 0.16%
Complete Rate 77.23% 72.66% 78.21% 80.47% 82.04% 79.78%
Personal Rate 71.18% 63.97% 73.77% 73.99% 75.98% 73.46%
4.3 Characteristics of the Dataset
Our LLM generator successfully identifies purchase reason in 96.4% of reviews. This includes 61.8% of reviews with explicitly stated reasons and 70.9% where reasons are inferred from the review and the product information. Our LLM identifies post-purchase experience in 88.2% of reviews, slightly lower than purchase reason. This is because we ask the LLM only to identify explicitly described experience, without making any inference. Table 4 shows explanations for two reviews in our dataset.
Overall 70.1% of purchase reasons and 71.18% of post-purchase experience are rated as personal. Among these, reviews for “Books” have the highest proportion of personalized purchase reason (78.89%) but the lowest proportion of personalized post-purchase experience. This is likely because buyers of books often start with a personal connection – why they chose the book, and then the focus shifts to an objective description of the book itself after reading, making the post-purchase experience less personalized.
We further characterize our dataset from linguistic aspects (detailed in Table 5). On average, purchase reasons are more concise (11.39 words and 10.16 words for explicit and implicit reason) than post-purchase experience (22.19 words). Both of them are much shorter than the original reviews (60.6 words). Since explanations are short, we use a very simple type to token ratio method (i.e., unique word count/total word count) to measure its lexical diversity  Templin (1957). We find that our generated explanations have a very high lexical diversity (close to 1) and rarely use duplicate tokens, while product information is the most likely to use duplicate tokens to advocate product features. Finally, we assess the dataset’s readability using the Flesch-Kincaid Grade Level metric Flesch (1948), a standard measurement in NLP research. Higher scores indicate greater readability. Our results align with expectations: product information and implicit purchase reasons (inferred from product information), are the most difficult to read. Conversely, user reviews and explicit purchase reasons are generally easier to understand.

### v2

@@ Abstract
In our work, we propose a novel purchase reason explanation task. To this end, we introduce an LLM-based approach to generate a dataset that consists of textual explanations of why real users make certain purchase decisions. We induce LLMs to explicitly distinguish between the reasons behind purchasing a product and the experience after the purchase in a user review. An automated, LLM-driven evaluation, as well as a small scale human evaluation, confirms the effectiveness of our approach to obtaining high-quality, personalized explanations. We benchmark this dataset on two personalized explanation generation tasks. We release code and prompts to spur further research [^1] The URL will be provided after the anonymity period..
@@
Tao Chen1  Siqi Zuo1  Cheng Li1  Mingyang Zhang1 Qiaozhu Mei2thanks: The work was done as a visiting researcher at Google DeepMind.Michael Bendersky1 1Google DeepMind  2 University of Michigan 1{taochen,siqiz,chgli,mingyang,bemike}@google.com 2qmei@umich.edu
@@ 1 Introduction
Providing user-understandable explanations to justify recommendations could enhance the effectiveness, persuasiveness, and user satisfaction of recommendation Zhang et al. (2020). This has been confirmed by both user studies Zanker (2012) and many real products such as Microsoft Office 365 Xu et al. (2020), JD.com (an e-commerce website) Zhang et al. (2014a) and Spotify McInerney et al. (2018). Explanations can be presented in various styles, such as a piece of text, a relevant user or item, a radar chart, an image or a set of reasoning rules Tintarev and Masthoff (2015). With the advancement of natural language generation techniques, a short narrative (e.g., a sentence) becomes one of the most popular explanation style, which is the focus of our work.
To pave the way of building a more effective recommendation system, we need datasets that explain the reasons behind user choices. Existing works leverage user reviews to mine explanations. One work Li et al. (2021a) extracts the most commonly occurred near-duplicate sentences across reviews, resulting in short and generic comments about the items. For instance, “Excellent movie” is the top 1 extracted explanation in Amazon Movies & TV review dataset. Others extract review sentences Geng et al. (2022) or segments Ni et al. (2019) that mention one or more pre-selected item features/aspects. An example explanation is “The quality of the material is great” where “quality” and “material” are two features of the item Ni et al. (2019).
@@
In this work, we introduce a novel purchase reason explanation task. Our approach leverages a large language model (LLM) to generate a dataset from user reviews, capturing purchase reasons and in the meantime, generating a highly relatable, personalized post-purchase experience. We propose three dimensions to measure the data quality and validate the effectiveness of our approach through automated LLM evaluator and small-scale human feedback. The resulting dataset is a high-quality, personalized set of explanations. With this dataset, we can develop models to generate explanations from two aspects, the relevance to the user need (i.e., purchase reason) and the preference for the particular item (i.e., post-purchase experience). We benchmark the dataset against these two generation tasks with various LLMs and explore the effectiveness of different user and item representations.
@@
  • To spur further research, we release the dataset, code and all the prompts for generating and benchmarking the dataset.
@@ 2 Related Work
In this section, we first review prior work on dataset construction for personalized explanations and explanation generation methods in the context of explainable recommender systems. We then discuss the use of LLMs for text generation and evaluation.
@@ 2.1 Personalized Explanation Dataset
Constructing a high quality personalized explanation dataset is the key to build an explainable model for personalized recommendation. Colas et al. (2023) use product description as ground truth explanation, which is fact-grounded but not personalized. The majority of prior work extracts various information (e.g., aspect-specific sentences) from user reviews Chen et al. (2018); Li et al. (2017); Chen et al. (2019); Wang et al. (2018b); Li et al. (2020); Geng et al. (2022); Ni et al. (2019). They discard segments that are too personal (containing first-person or third-person pronouns) or too short/long. Li et al. (2021a) extract commonly occurred near-duplicate sentences as explanations, which are often short and generic (e.g., “Excellent movie”). The above approaches capture customer satisfaction rather than purchase motives because they do not distinguish pre- from post-purchase experiences.
Li et al. (2017) use Tips [^2] https://www.yelp-support.com/article/What-are-tips?l=en_US, a concise form of reviews, as justifications for choosing certain businesses. But Tips have low availability (e.g., Yelp mobile). To our knowledge, we are the first to extract purchase reasons directly from regular reviews for creating personalized explanation datasets.
@@ 2.2 Personalized Explanation Generation
Existing approaches, such as ranking-based Li et al. (2021a, 2023a) and template-based Zhang et al. (2014b); Wang et al. (2018a), tend to produce generic explanations with limited language flexibility and personalization.
Natural language generation (NLG) is widely used to generate free-text explanations. Early studies fine-tuned seq2seq models like LSTM Costa et al. (2018), GRU Li et al. (2017); Ni et al. (2019), Transformer Li et al. (2021b), T5 Liu et al. (2023) or GPT-2 Li et al. (2023b). The results were often generic and not fluent Liu et al. (2023). Recent work with powerful LLMs (e.g., ChatGPT Liu et al. (2023)) in zero or few-shot setups has significantly improved quality.
A common challenge for NLG-based methods is effectively representing users and items. Most studies use textual descriptions or IDs Liu et al. (2023); Li et al. (2023b), with the latter limiting generalization to new IDs.
A shared challenge of all these approaches is systematically evaluating generated personalized explanations, which we address by introducing a new dataset and benchmark.
2.3 LLMs for Text Generation and Evaluation
While LLMs demonstrate exceptional capabilities in generating text Zhao et al. (2023), their tend to hallucinate Zhang et al. (2023); Huang et al. (2023). LLMs have been widely used to evaluate the quality of generated text across various domains Zhao et al. (2023); Wang et al. (2023a); Kocmi and Federmann (2023); Wang et al. (2023b), often achieving state-of-the-art or competitive correlation with human judgments.
@@ 3 Generating the Explanation Dataset
As we discussed in Section 2.1, existing work extracting explanations from reviews suffers from two key limitations. First, they most often rely on sentiment, or attitude towards the item, which usually stems from user experience after purchasing the item. Consequently, it might not explain the decision to make the purchase. Second, existing methods often produce generic explanations (e.g.,  Li et al. (2021a) use frequently occurring phrases and  Ni et al. (2019) discard text segments that are personal), overlooking the personal nature of purchase motivations.
@@
To address these limitations, we explicitly distinguish the purchase reasons from post-purchase experience. These two concepts have strong roots in the business and marketing literature, where purchase reason is related to purchase intent (or intention) Chang and Wildt (1994) and purchase motive (or motivation) Sirgy (1985), while post-purchase experience is related to customer (or product) satisfaction Oliver and Swan (1989); Richins and Bloch (1991). In particular, purchase intent refers to the likelihood that a user will buy a product, which is influenced by a number of factors such as personal needs, customer preferences, brand equity, trust, perceived value of the product, etc. Cobb-Walgren et al. (1995); Bleize and Antheunis (2019) In our work, we do not further distinguish these factors and in general refer to them as purchase reasons. On the other hand, post-purchase experience encodes a variety of factors that influences post-purchase customer satisfaction, which is highly related to the user’s attitude towards the product or business and may further influence their future purchasing behavior Anderson et al. (1979). In our work, the practical difference between purchase reason and post-purchase experience lies in whether the conveyed information is before the user purchases the item or after. The former explains the user’s decision to purchase the item, and the latter explains the decision of rating the item.
@@
As LLMs have demonstrated strong performance in text generation, we propose to use LLMs to generate both types of explanations based on user reviews, rather than adopting the extractive methods proposed by prior work. We carefully devise strategies to improve generation quality and combat hallucinations. Moreover, we use LLMs as auto-raters to assess the generated explanations from multiple aspects and further improve data quality. Specifically, we demonstrate the effectiveness of our solution on the Amazon product review dataset  [^3] https://nijianmo.github.io/amazon/index.htmlNi et al. (2019) employing a diverse range of LLMs. These include two commercial models, Gemini Ultra Gemini Team (2023) and GPT-4 Turbo OpenAI (2023), as well as an open-source LLM Gemma-7B Gemma Team (2024). We detail our end-to-end solution below.
3.1 Generating Explanations using LLMs
Given a product and an associated user review, our goal is to use an LLM to generate this user’s purchase reason and post-purchase experience. Initially, we treated the two generation tasks separately but encountered two issues. First, LLMs struggle to distinguish between the two tasks, leading to overlapping or miscategorized explanations. Second, when reviews lack explicit purchase reasons, the model either infers them from product information or hallucinates. We address these issues with strategies detailed below, with our final prompt in Appendix (Table 12).
1. We instruct the model to perform the two generation tasks simultaneously in one prompt, enforcing a clear distinction between purchase reasons and post-purchase experiences.
2. To reduce hallucination, we separate purchase reasons into explicit (directly stated in the review) and implicit (inferred from the review and product information). The model can leave either blank if information is not found.
@@ 3.1 Generating Explanations using LLMs
3. We require the model to provide supporting evidence for its generated purchase reasons. This further helps combat hallucination.
4. We incorporate two examples for in-context learning: one demonstrating both purchase reasons and post-purchase experiences, and the other featuring a concise review containing only post-purchase experience. Our findings suggest that these examples effectively mitigate the generation of false purchase reasons and enhance the completeness of post-purchase experiences, especially for short reviews.
3.2 Rating Explanations using LLMs
@@
Prior work has demonstrated a great success of automatic evaluation of text generation with LLMs (detailed in Section 2.3). Inspired by this, we explore the usage of LLMs in assessing the quality of our explanation dataset. In particular, we propose to measure the quality of two types of explanations (i.e., purchase reasons and post-purchase experience) from the following three dimensions:
@@
Completeness: if the explanation covers all relevant aspects present in the review and the product.
Type Match: if the model correctly identifies the explanation type. We observe that the model sometimes confuse post-purchase experience as purchase reason, but rarely the opposite.
We conduct two separate evaluations for the two types of explanations (see Table 13 and Table 14 in Appendix for the specific prompts). We further leverage the three evaluation results to filter out noises, aiming at improving the dataset quality.
@@ 4 Dataset Evaluation and Analysis
We use the Amazon product review 5-core dataset for experiments.The full 5-core dataset consists of reviews from all users and items that have at least 5 reviews, resulting in 75 million reviews in total. Considering the cost of using LLMs, we randomly sample 10K reviews for experiments. We apply the LLM generator to construct the explanation dataset and utilize the LLM auto-rater to judge the data quality. For purchase reason, we combine the explicit and implicit reasons into one for evaluation, analysis and later benchmark (Sec. 5). In the following, we first validate the effectiveness of our LLM auto-rater and then describe the quality and characteristics of the generated explanation dataset. We choose Gemini Ultra as the primary LLM for both generator and auto-rater, but also assess the generalizability of our approach using GPT-4 Turbo and the more compact, open-source Gemma-7B.
@@ 4.1 Effectiveness of LLM Auto-rater
To evaluate our LLM auto-rater, we conduct a small scale human annotations following the same guideline used by the auto-rater. In the pilot study (20 reviews, 4 annotators), agreement was high: perfect consensus on hallucination, and 90% agreement on completeness and type match. Disagreements were resolved through discussion. Subsequently, a larger set of 100 randomly sampled reviews are annotated (one annotator per review).
Table 2 shows that the LLM auto-rater is highly effective, achieving perfect correlation with human judgment for hallucination and type match. Agreement on completeness reached 92% for purchase reasons and 95% for post-purchase experiences. Analysis suggests the auto-rater is stricter than humans for post-purchase experience completeness but more lenient for purchase reason completeness, indicating that identifying purchase reasons may be a more challenging task.
Table 2: The agreement rate between the LLM auto-rater and human annotator on 100 sampled reviews.
Hallucination Completeness Type Match
Reason 100% 92% 100%
Experience 100% 95%
@@ 4.2 Quality of Generated Explanations
Table 3: The auto-rater results on 10K dataset. Fashion denotes Clothing, Shoes & Jewelry.
All Fashion Sports & Outdoors Books Electronics Home & Kitchen
@@
Hallucination Rate 0.91% 0.55% 1.24% 1.19% 1.10% 1.01%
Completeness Rate 89.41% 86.85% 91.34% 90.02% 90.11% 88.10%
@@
Hallucination Rate 0.86% 0.55% 0.62% 0.81% 1.65% 0.82%
Completeness Rate 95.88% 96.01% 97.06% 95.52% 96.37% 95.79%
Type Match Rate 99.11% 99.24% 98.92% 99.08% 98.90% 98.81%
Table 3 shows the evaluation of the generated explanations using the LLM auto-rater. Top product categories are also shown for understanding quality variations. The LLM generator achieves a low hallucination rate (0.91% for purchase reasons, 0.86% for post-purchase experiences) and rarely confuses the two generation tasks (0.89%).
The completeness rate is lower than other two dimensions (89.41% for purchase reasons, 95.88% for post-purchase experience). As discussed in Sec. 4.1, the strict auto-rater considers subtle aspects for the completeness evaluation, suggesting that the actual rate might be higher.
Within the five product categories, “Sports & Outdoors” has the highest completeness rate for both purchase reason (91.34%) and post-purchase experience (97.06%). “Clothing, Shoes & Jewelry” (86.85%) has the lowest rate for purchase reasons, while “Books” (95.52%) has the lowest for post-purchase experiences. This discrepancy might be due to varying purchase and review behaviors across categories. Unlike “Sports & Outdoors” reviews, where buyers often state their motivations explicitly, reviews for clothing, shoes, or jewelry may require the LLM to infer motivations from subtle cues, potentially leading to omissions. Conversely, the descriptive language in book reviews, compared to the brevity of “Sports & Outdoors” reviews, poses a challenge for the LLM to capture the nuances of reader sentiment.
4.3 Characteristics and Differences to Other Datasets
Our LLM generator successfully identifies purchase reason in 93.94% of reviews. This includes 36.8% of reviews with explicitly stated reasons and 84.2% where reasons are inferred from the review and the product information. The majority of reviews lacking an identified purchase reason are short (55% are shorter than 20 tokens), suggesting that users may not mention or provide any clues regarding their motivation for buying in these cases. Our LLM identifies post-purchase experience in nearly all reviews (99.93%). This is because users typically express their satisfaction or dissatisfaction in the reviews, even in very short ones.
Table 4 shows explanations for two reviews in our dataset. To further differentiate our dataset, we showcase a few examples from two previous explanation datasets derived from Amazon reviews: EXTRA Li et al. (2021a), which utilizes commonly occurring near-duplicate phrases as explanations, and P5 Geng et al. (2022), which extracts sentences containing product features as explanations, alongside our own explanations at Table 19 and Table 20 in the Appendix.
We further characterize our dataset from linguistic aspects (detailed in Table 5). On average, generated purchase reasons (10.63 words) are shorter than post-purchase experiences (23.43 words), and both are significantly shorter than the original reviews (63.79 words). As expected, explanations in the EXTRA dataset are very short (4.96 words), while P5 explanations are longer (25.33 words). The type-to-token ratio (i.e., unique word count/total word count) Templin (1957) shows that our generated explanations have significantly higher vocabulary diversity (0.140 for purchase reason and 0.048 for post-purchase experience) compared to the EXTRA (0.006) and P5 (0.014) datasets.